# Controlled Retrieval-augmented Context Evaluation for Long-form RAG

### v1

@@ 4.2 Controlled Empirical Experiments
What are the reference performance bounds of retrieval context and final RAG result?
@@
Unsurprisingly, we observe the lowest coverage of RAG result without retrieval (#1), confirming that parametric knowledge in the LLM alone is insufficient to achieve high performance. This condition serves as the empirical lower bound of RAG. In contrast, the oracle result using the human-written summary (#2) achieves highest coverage with answering over 90% of sub-questions. It implies that generated sub-questions are answerable and validate the framework’s ability to capture completeness. The RAG result with oracle retrieval context (#3) yields decent coverage of 64.6 and 61.8, outperforming other empirical methods in subsequent blocks in the table. This demonstrates an empirical upper bound for RAG’s retrieval, grounded in an oracle retrieval context $Z^{*}$. Overall, CRUX provides robust bounds for reference, enabling more diagnostic evaluation of RAG’s retrieval regardless of the generator’s effects.
@@
To investigate this, we evaluate a range of empirical retrieval contexts from various cascaded retrieval pipelines. As reported in Table 1, each pipeline is evaluated with both the quality of intermediate retrieval context $Z$ and the final RAG result $y$ (the gray columns).
@@
Notably, on evaluation results of DUC, we observe pointwise re-ranking models have robust gains on final RAG result coverage only when used with weaker initial retrieval (e.g., LR + miniLM, 35.8 $\rightarrow$ 38.4). However, they degrade when adopting stronger initial retrieval (e.g., LSR + miniLM, 41.0$\rightarrow$ 39.2). Such patterns are also shown on intermediate retrieval context performance, demonstrating CRUX’s evaluation capability for retrieval context.
In contrast, more effective re-ranking consistently enhances overall performance, with visible performance gains in both intermediate and final results. For example, RankFirst Reddy et al. (2024) and SetwiseFlanT5 Zhuang et al. (2023), particularly outperform all the other empirical pipelines (conditions marked in bold). Yet, they still have a large gap compared to the oracle retrieval (#3), implying that existing ranking models are not explicitly optimized for coverage of long-form RAG results.
Can intermediate retrieval context performance extrapolate the final RAG result performance?
Finally, to highlight the advantage of retrieval context evaluation, we compute the ranking correlation in terms of Kendall $\tau$ between final result coverage/density (i.e., $Cov(y)$/$Den(y)$) and the intermediate coverage, ranked coverage and density.
We find ranking correlation strengths of approximately 0.7 to 0.8 on both evaluation sets at the last row in Table 1, demonstrating the strong alignment between retrieval context and RAG result. This suggests that our framework can be a promising surrogate retrieval evaluation for extrapolating long-form RAG results.
@@ 4.3 Metric Alignment Analysis
Figure 4: Coverage of RAG results for 10 CRUX-DUC queries ($x$-axis) under three retrieval contexts ($y$-axis). Each subplot shows LLM-judged coverage (line) and human judgments (markers); bars indicate the annotators’ average. The Pearson correlations $\rho$ are computed between the LLM and each annotator’s coverage.
@@
We conduct human judgment on 10 randomly selected open-ended queries from CRUX-DUC. We design two reading comprehension tasks: [^8] Appendix A.2 details the annotation tasks (e.g., process, interface design and annotator, etc.). $\mathcal{T}_{1}$: Long-form RAG result coverage judgement, and $\mathcal{T}_{2}$: Rubric-based passage judgement. $\mathcal{T}_{1}$ investigates how well LLM-judged coverage align with human’s. We collect binary answerability annotations for all enumerated result sub-question pairs $\{(y,q_{1}),...,(y,q_{n})\}$ and compute the corresponding coverage of final RAG result $Cov(y)$.
We evaluate RAG results across three retrieval contexts $Z$: Oracle, BM25 and DR+RankFirst, as shown in the subplots in Figure 4. With the total of 30 human-judged coverage, we compute the Spearman correlation between them and LLM, obtaining high alignment ($\rho\geq 0.8)$, and a moderate inter-annotator agreement (Fleiss’ $\kappa=0.52$). We also found that the controlled oracle retrieval $Z^{*}$ has significantly better coverage from human judgements, confirming the reliability of upper bound, while the other retrieval context are fluctuate among queries.
@@
Figure 5: Kendall $\tau$ rank correlations between evaluation metrics on CRUX-DUC, using 48 random sampled retrieval contexts $Z$. Metrics includes intermediate and final coverage, and other relevance-based metrics.
@@ 4.4 Configuration Analysis
We first adjust the higher answerability threshold ($\eta=5$ in Eq. (2)). Our analysis is conducted on CRUX-DUC evaluation set using the same empirical retrieval pipelines. In Table 2, we observe the higher threshold leads to lower coverage in both intermediate and final results, $Cov(Z)$ and $Cov(y)$. While setting threshold as 3 demonstrates slightly larger variance (±$\pm 3$) across retrieval pipelines, which is more discriminative and desirable. Similarly, we compute the ranking correlations under two thresholds and justify that $\eta=3$ achieves better alignment; we thereby set it as default throughout this study.
@@
Table 3: Kendall $\tau$ rank correlations between intermediate retrieval context and final result evaluation, under different size of retrieval context and datasets. The columns 2 to 6 compare with final coverage $Cov(y)$ and the last column compares final density $Den(y)$.
@@ Acknowledgments
We acknowledge the Dutch Research Council (NWO) in The Netherlands for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the ‘Computing Time on National Computer Facilities’ call.
@@ A.1 Empirical Evaluation
We adopt larger Llama models Meta (2024), Llama-3.1-70B-Instruct, to generate the CRUX evaluation datasets: CRUX-DUC and CRUX-Multi-News (test split). For training data generation using the Multi-News train split, we employ Llama-3.1-8B-Instruct due to the high computational cost of large-scale generation. Generation is performed under two different settings. For text generation (e.g., queries, passages, and questions), we use a temperature of 0.7 and top-$p$ of 0.95. For judgement generation (i.e., graded ratings for answerability), we follow Thomas et al. (2024) and use a temperature of 0.0 and top-$p$ of 1.0. To accelerate inference, we leverage vLLM Kwon et al. (2023). The entire data generation process is conducted on 4 AMD MI200X GPUs and takes approximately 14 days.
@@
Sub-questions Generation Instruction: Write {$n$} diverse questions that can reveal the information contained in the given document. Each question should be self-contained and have the necessary context. Write the question within ‘<q>’ and ‘</q>’ tags. Document: {$c^{*}$} Questions: <q>
@@
Graded Rating Generation Instruction: Determine whether the question can be answered based on the provided context? Rate the context with on a scale from 0 to 5 according to the guideline below. Do not write anything except the rating. Guideline: 5: The context is highly relevant, complete, and accurate. 4: The context is mostly relevant and complete but may have minor gaps or inaccuracies. 3: The context is partially relevant and complete, with noticeable gaps or inaccuracies. 2: The context has limited relevance and completeness, with significant gaps or inaccuracies. 1: The context is minimally relevant or complete, with substantial shortcomings. 0: The context is not relevant or complete at all. Question: {$q$} Context: {$c$} Rating:
@@
[…]
(#1) What was the initial reaction of the audience when Colin Yost started dancing during his commencement speech?
(#2) How did Colin Yost prepare his classmates for the surprise dance routine?
(#3) What song did Colin Yost choose for the flash mob dance routine?
(#4) What was the main theme of Colin Yost’s commencement speech?
(#5) What did Colin Yost plan to study in college?
(#6) What was the audience’s reaction to the flash mob dance routine?
(#7) How did Colin Yost convince the school administration to allow the flash mob dance routine during the graduation ceremony?
(#8) What college will Colin Yost be attending in the fall?
2. How many students from Portsmouth High School’s senior class participated in the choreographed dance celebration? 8. Did Colin Yost have any prior dance training before the graduation ceremony?
[…]
Answerability (3/10) : [0, 0, 5, 5, 0, 0, 0, 0, 5, 0] –> {#2, #3, #7}
[…]
Answerability (3/10) : [5, 0, 0, 0, 5, 0, 5, 0, 0, 0] –> {#1, #4, #6}
[…]
Answerability (4/8): [5 0 0 0 5 5 0 5] –> {#1, #5, #6, #8}
@@ A.2 Human Evaluation
We include 30 machine-generated RAG results (reports), with each result containing 15 sub-questions to be labeled as either answerable or unanswerable. The guideline is reported in Figure 10. The 30 reports are from three types of retrieval contexts: Oracle, BM25, and DR+RankFirst (10 each), to ensure a balanced distribution across retrieval settings. The human coverage reported in Figure 4 is calculated in line with LLM judgement using the same set of answerable sub-questions (see Sec. 3.3).
Annotation task 2–passage-level judgement with rubric-based graded rating.
In $\mathcal{T}_{2}$, we randomly select oracle relevant passages and ask annotators to label graded ratings from 0 to 5 for two random sub-questions, simulating the LLM-based judgement using the prompt shown in Figure 8. We collected 226 human ratings (ground truth) and compared them to LLM predictions. We observe precision above 0.6 for both answerable ($\eta\geq 3$) and unanswerable ($\eta<3$) cases. While recall is high for unanswerable questions, it drops to 0.4 for answerable ones. This indicates the LLM tends to make conservative predictions, underestimating answerable content. A key challenge for improving CRUX is generating sub-questions that are both more discriminative and better aligned with human perception.
@@ A.3 Case Study
Annotation Task 1a: Answerability Judgement Your first step is to evaluate whether each of the 15 questions (Q-0 to Q-14) is answerable based solely on the machine-generated report. Carefully read the entire report before starting the questions (the open-ended query is just for you reference). Click the corresponding button (e.g. Q-0, Q-1, etc.) to view the question. Decide if the report contains enough information to answer the question. If the report provides enough information to answer the question, select "1 (Answerable)". If the report does not provide any information, select "0 (Unanswerable)". (Note) Your judgment should be based on whether the information is present. You do not need to verify external truth.

### v2

@@ 4.2 Controlled Empirical Experiments
What are the reference performance bounds of the retrieval context and final RAG result?
@@
Unsurprisingly, we observe the lowest coverage for RAG result without retrieval (#1), confirming that parametric knowledge in the LLM alone is insufficient to achieve high performance. This condition serves as the empirical lower bound of RAG. In contrast, the oracle result using the human-written summary (#2) achieves the highest coverage by answering over 90% of the sub-questions. This implies that the generated sub-questions are answerable and validate the framework’s ability to capture completeness. The RAG result with the oracle retrieval context (#3) yields decent coverage of 64.6 and 61.8, outperforming other empirical methods in subsequent blocks in the table. This demonstrates an empirical upper bound for RAG’s retrieval, grounded in an oracle retrieval context $Z^{*}$. Overall, CRUX provides robust bounds for reference, enabling more diagnostic evaluation of RAG’s retrieval regardless of the generator.
@@
To investigate this, we evaluate a range of empirical retrieval contexts from various cascaded retrieval pipelines. As reported in Table 1, each pipeline is evaluated with both the quality of the intermediate retrieval context $Z$ and the final RAG result $y$ (the gray columns).
@@
[…]
In contrast, more effective re-ranking methods consistently enhances overall performance, with visible performance gains in both intermediate and final results. For example, RankFirst Reddy et al. (2024) and SetwiseFlanT5 Zhuang et al. (2023), particularly outperform all the other empirical pipelines (conditions marked in bold). Yet, they still have a large gap compared to the oracle retrieval (#3), implying that existing ranking models are not explicitly optimized for coverage of long-form RAG results.
Can intermediate retrieval context performance extrapolate to final RAG result performance?
Finally, to highlight the advantage of retrieval context evaluation, we compute the ranking correlation in terms of Kendall’s $\tau$ between the final result coverage/density (i.e., Cov(y)Cov(y)/$Den(y)$) and the intermediate coverage, ranked coverage and density.
We find ranking correlation strengths of approximately 0.7 to 0.8 on both evaluation sets at the last row in Table 1, demonstrating the strong alignment between the retrieval context and RAG results. This suggests that our framework can be a promising surrogate retrieval evaluation for extrapolating long-form RAG results.
@@ 4.3 Metric Alignment Analysis
Figure 4: Coverage of RAG results for 10 CRUX-DUC queries ($x$-axis) under three retrieval contexts ($y$-axis). Each subplot shows LLM-judged coverage (line) and human judgments (markers); bars indicate the annotators’ average. The Spearman correlations $\rho$ are computed between the LLM and each annotator’s coverage.
@@
We conduct human judgment on 10 randomly selected open-ended queries from CRUX-DUC. We design two reading comprehension tasks: [^8] Appendix A.2 describes the annotation tasks in detail. $\mathcal{T}_{1}$: Long-form RAG result coverage judgment, and $\mathcal{T}_{2}$: Rubric-based passage judgment. $\mathcal{T}_{1}$ investigates how well LLM-judged coverage align with human’s. We collect binary answerability annotations for all enumerated result sub-question pairs $\{(y,q_{1}),...,(y,q_{n})\}$ and compute the corresponding RAG result’s coverage $Cov(y)$.
We evaluate RAG results across three retrieval contexts $Z$: Oracle, BM25 and DR+RankFirst, as shown in the subplots in Figure 4. With the total of 30 human-judged coverage, we compute the Spearman correlation between them and LLM, obtaining high alignment ($\rho\geq 0.8)$, and a moderate inter-annotator agreement (Fleiss’ $\kappa=0.52$). We also found that the controlled oracle retrieval $Z^{*}$ has significantly better coverage from human judgments, confirming the reliability of upper bound, while the other retrieval context fluctuate among queries.
@@
Figure 5: Kendall $\tau$ rank correlations between evaluation metrics on CRUX-DUC, using 48 random sampled retrieval contexts $Z$. Metrics include intermediate and final coverage, and other relevance-based metrics.
@@ 4.4 Configuration Analysis
We first adjust the higher answerability threshold ($\eta=5$ in Eq. (2)). Our analysis is conducted on CRUX-DUC evaluation set using the same empirical retrieval pipelines. In Table 2, we observe the higher threshold leads to lower coverage in both intermediate and final results, $Cov(Z)$ and $Cov(y)$. While setting threshold as 3 demonstrates slightly larger variance ($\pm 3$) across retrieval pipelines, which is more discriminative and desirable. Similarly, we compute the ranking correlations under two thresholds and justify that $\eta=3$ achieves better alignment; we thereby set it as default throughout this study.
@@
Table 3: Kendall $\tau$ rank correlations between the intermediate retrieval context and final result evaluation, with retrieval context sizes and datasets. The columns 2 to 6 compare with final coverage $Cov(y)$ and the last column compares final density $Den(y)$.
@@ Acknowledgments
This research was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl, project VI.Vidi.223.166 of the NWO Talent Programme which is (partly) financed by the Dutch Research Council (NWO), projects NWA.1389.20.183 and KICH3.LTP.20.006, and the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE). We acknowledge the Dutch Research Council for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through project number NWO-2024.050. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of their respective employers, funders and/or granting authorities.
@@ A.1 Empirical Evaluation
We adopt larger Llama models Meta (2024), Llama-3.1-70B-Instruct, to generate the CRUX evaluation datasets: CRUX-DUC and CRUX-Multi-News (test split). For training data generation using the Multi-News train split, we employ Llama-3.1-8B-Instruct due to the high computational cost of large-scale generation. Generation is performed under two different settings. For text generation (e.g., queries, passages, and questions), we use a temperature of 0.7 and top-$p$ of 0.95. For judgment generation (i.e., graded ratings for answerability), we follow Thomas et al. (2024) and use a temperature of 0.0 and top-$p$ of 1.0. To accelerate inference, we leverage vLLM Kwon et al. (2023). The entire data generation process is conducted on 4 AMD MI200X GPUs and takes approximately 14 days.
@@
Sub-questions Generation Instruction: Write {nn} diverse questions that can reveal the information contained in the given document. Each question should be self-contained and have the necessary context. Write the question within ‘<q>’ and ‘</q>’ tags. Document: $c^{*}$} Questions: <q>
@@
Graded Rating Generation Instruction: Determine whether the question can be answered based on the provided context? Rate the context with on a scale from 0 to 5 according to the guideline below. Do not write anything except the rating. Guideline: 5: The context is highly relevant, complete, and accurate. 4: The context is mostly relevant and complete but may have minor gaps or inaccuracies. 3: The context is partially relevant and complete, with noticeable gaps or inaccuracies. 2: The context has limited relevance and completeness, with significant gaps or inaccuracies. 1: The context is minimally relevant or complete, with substantial shortcomings. 0: The context is not relevant or complete at all. Question: {qq} Context: {cc} Rating:
@@
Open-ended Query. Research the graduation ceremony of Portsmouth High School in New Hampshire and write a report on the activities that took place during the event. Include details on the valedictorian’s speech and the surprise dance routine performed by the graduating class.
Sub-questions. (2 questions are filtered by Oracle Passages) (#1) What was the initial reaction of the audience when Colin Yost started dancing during his commencement speech? (#2) How did Colin Yost prepare his classmates for the surprise dance routine? (#3) What song did Colin Yost choose for the flash mob dance routine? (#4) What was the main theme of Colin Yost’s commencement speech? (#5) What did Colin Yost plan to study in college? (#6) What was the audience’s reaction to the flash mob dance routine? (#7) How did Colin Yost convince the school administration to allow the flash mob dance routine during the graduation ceremony? (#8) What college will Colin Yost be attending in the fall? 2. How many students from Portsmouth High School’s senior class participated in the choreographed dance celebration? 8. Did Colin Yost have any prior dance training before the graduation ceremony?
Oracle Passage. #1. Colin Yost, the valedictorian at Portsmouth High School in Portsmouth, New Hampshire, delivered an unforgettable commencement speech that ended with a surprise dance routine to Taylor Swift’s S̈hake It Off.Ḧe had been planning this moment for some time, inspired by his desire to do a flash mob and showcase his class’s cohesion. Yost worked with a few friends to choreograph the dance and shared an instructional video with the class on YouTube. The administration was on board with the plan, allowing the seniors to use five graduation rehearsals to perfect the routine. Answerability (3/10) : [0, 0, 5, 5, 0, 0, 0, 0, 5, 0] –> {#2, #3, #7}
Oracle Passage #2. As Yost began his speech, he emphasized the importance of embracing one’s inner nerd and striving for perfection in anything one is passionate about. He then ended his speech with the iconic line äll you have to do is shake it off,b̈efore breaking into dance. The initial reaction was mixed, with some parents laughing and others looking confused. However, as the front row joined in, followed by another row, the energy shifted, and the audience was soon filled with laughter and tears. Answerability (3/10) : [5, 0, 0, 0, 5, 0, 5, 0, 0, 0] –> {#1, #4, #6}
Oracle Passage #3. Yost’s creative and entertaining approach to his commencement speech has gained attention, especially during a season when many notable figures, including President Obama and Stephen Colbert, have been delivering inspiring speeches. Yost’s message of embracing individuality and having fun was well-received by his classmates and their families. As he prepares to attend Princeton in the fall, where he plans to major in chemical and biological engineering, Yost’s unique approach to his commencement speech will undoubtedly be remembered. Answerability (2/10) : [0, 0, 0, 0, 5, 5, 0, 0, 0, 5] –> {(#4), #5, #8}
Oracle Result (human-written summary). Parents who thought they were going to have to sit through a boring graduation in a stuffy gym got anything but at Portsmouth High School on Friday. Colin Yost, the valedictorian for the New Hampshire school’s senior class, decided he wanted to shake things up—and off—during his commencement speech, so after his words of inspiration, he stepped out from behind the podium and (#3:) started dancing, by himself, to Taylor Swift’s "Shake It Off," eliciting laughter and some öh gosh, what is he doing?r̈eactions, MTV reports. Soon, however, his intentions were made clear as the rest of his graduating class (more than 230 in all) stood up and joined Colin in a choreographed celebration of the end of their high school career. While Colin’s flash mob surprised the audience, it was far from spontaneous. (#2:) The senior posted a video tutorial on YouTube for his classmates to study and (#7:) cajoled the school’s administration beforehand into letting him use five graduation rehearsals to get the moves down just right, MTV notes. Äs we practiced, the energy was just building and everyone was feeling how great it was to work together and send this positive message,ḧe tells the station. He adds that the song-and-dance show played perfectly into what he had talked about in his speech on embracing your inner nerd, the Portsmouth Herald notes. But despite the Taylor-made two-stepping’s success, we probably won’t be seeing Colin—who admits he’s never taken a dance lesson—on So You Think You Can Dance: He’s headed to Princeton to study chemical and biological engineering, per MTV. (Hopefully no one got arrested for cheering.) Answerability (4/8): [5 0 0 0 5 5 0 5] –> {#1, #5, #6, #8}
@@ A.2 Human Evaluation
We include 30 machine-generated RAG results (reports), with each result containing 15 sub-questions to be labeled as either answerable or unanswerable. The guideline is reported in Figure 10. The 30 reports are from three types of retrieval contexts: Oracle, BM25, and DR+RankFirst (10 each), to ensure a balanced distribution across retrieval settings. The human coverage reported in Figure 4 is calculated in line with LLM judgment using the same set of answerable sub-questions (see Sec. 3.3).
Annotation task 2–passage-level judgment with rubric-based graded rating.
In $\mathcal{T}_{2}$, we randomly select oracle relevant passages and ask annotators to label graded ratings from 0 to 5 for two random sub-questions, simulating the LLM-based judgment using the prompt shown in Figure 8. We collected 226 human ratings (ground truth) and compared them to LLM predictions. We observe precision above 0.6 for both answerable ($\eta\geq 3$) and unanswerable ($\eta<3$) cases. While recall is high for unanswerable questions, it drops to 0.4 for answerable ones. This indicates the LLM tends to make conservative predictions, underestimating answerable content. A key challenge for improving CRUX is generating sub-questions that are both more discriminative and better aligned with human perception.
@@ A.3 Case Study
Annotation Task 1a: Answerability Judgment Your first step is to evaluate whether each of the 15 questions (Q-0 to Q-14) is answerable based solely on the machine-generated report. Carefully read the entire report before starting the questions (the open-ended query is just for you reference). Click the corresponding button (e.g. Q-0, Q-1, etc.) to view the question. Decide if the report contains enough information to answer the question. If the report provides enough information to answer the question, select "1 (Answerable)". If the report does not provide any information, select "0 (Unanswerable)". (Note) Your judgment should be based on whether the information is present. You do not need to verify external truth.